OpenStudy (anonymous):

Multiply (y+7)(y-1) Using the FOIL method.

OpenStudy (anonymous):

Try it yourself first. What do you get?

OpenStudy (anonymous):

Its not for me its for a girl tht needs help with her homework. I honestly do not remember how to do this.

OpenStudy (anonymous):

FOIL here means ?

OpenStudy (anonymous):

FOIL = First Outer Inner Last It's just distributive property.

OpenStudy (anonymous):

do y*y = y^2 then y * -1 = -y then 7 * y = 7y then 7 * -1 = -7 so you have y^2 - y +7y - 7 so its, y^2 + 6y - 7 is your answer

OpenStudy (anonymous):

I just did it for her, review my steps then you will see